The two most prominent wavelengths in the light emitted by a hydrogen discharge lamp are 656 nm (red) and 486 nm (blue). Light from a hydrogen lamp illuminates a diffraction grating with 550 lines / mm, and the light is observed on a screen 1.2 m behind the grating.

You may want to review (Page 940).

For general problem-solving tips and strategies for this topic, you may want to view a Video Tutor Solution of Diffraction grating.


Part A

What is the distance between the first-order red and blue fringes?
Answer #1

Here,

distance from screen , D = 1.2 m

slit separation , d = 1/550 mm = 10^-3/550 m

Now, for the distance between first order red and blue fringe

y = D * difference in wavelength/d

y = 1.2 * (656 - 486) *10^-9/(10^-3/550)

y = 0.1122 m

the distance between the first order red and blue fringes is 0.1122 m